OpenStudy (anonymous):

Hey Can Someone Please Help Me? /.\ If So Thank You <3 (i'll fan) Solve for x: 2 over 5 (x - 4) = 2x.

jigglypuff314 (jigglypuff314):

to get rid of denominators you can multiply by 5 so 2(x-4) = 5(2x) distribute/multiply what would you get? :)

OpenStudy (anonymous):

i dont understand distrubuting. /.\

jigglypuff314 (jigglypuff314):

distributing is like a(b-c) = ab - ac for example 3(x-5) = 3x - 15 so now try-> 2(x-4) = ?

OpenStudy (anonymous):

2x-8?

jigglypuff314 (jigglypuff314):

yes! :)

OpenStudy (anonymous):

okay! so now what >.< lol

jigglypuff314 (jigglypuff314):

so from 2(x-4) = 5(2x) you can get 2x - 8 = 10x from there you can subtract 2x from both sides :)

OpenStudy (anonymous):

8=5x?

jigglypuff314 (jigglypuff314):

not quite :) you would subtract, not divide so 2x - 8 = 10x -2x -2x

OpenStudy (anonymous):

ohhh so 8=8x?

jigglypuff314 (jigglypuff314):

yep :) then you would divide both sides by 8

OpenStudy (anonymous):

thank you!<3

jigglypuff314 (jigglypuff314):

you're welcome :)